Two angles of a polygon are right angle and the remaining are 120 each.find the no. Of sides in it

Answer by Alan3354(69443)

Two angles of a polygon are right angle and the remaining are 120 each.find the no. Of sides in it
----------
Sum of angles = 2*90 + (n-2)*120
Sum = 180*(n-2) = 180 + n*120 - 240
180n - 360 = 120n - 60
60n = 300
n = 5